147N+42He=178O+?? a. 10n b. 11H c. 42He d. 0-1e

147N+42He=178O+??

a. 10n

b. 11H

c. 42He

d. 0-1e
Answer #1

11H

mass number on left = mass number on right

14 + 4 = 17 + A

A = 1

atomic number of left = atoimic number on right

7 + 2 = 8 + Z

Z = 1

so 11H